OpenStudy (anonymous):

Some steps to rewrite the expression x3 - 9x + x2 - 9 as a product of three factors are shown below: Step 1: x3 - 9x + x2 - 9 Step 2: x3 + x2 - 9x - 9 Step 3: x2(x + 1) - 9(x + 1) Which of the following best shows the next two steps to rewrite the expression? Step 4: (x2 + 9)(x + 1); Step 5: (x + 3)(x + 3)(x + 1) Step 4: (x2 - 9)(x + 1); Step 5: (x + 3)(x + 3)(x + 1) Step 4: (x2 + 9)(x + 1); Step 5: (x - 3)(x + 3)(x + 1) Step 4: (x2 - 9)(x + 1); Step 5: (x - 3)(x + 3)(x + 1)

OpenStudy (anonymous):

@chrissyC. @cookielate @ooxBRITTANYxoo

OpenStudy (anonymous):

anyone?

OpenStudy (anonymous):

oh my

OpenStudy (anonymous):

i have nooooo clue

OpenStudy (anonymous):

Ok well when you originally solve an expression how would you do it?

OpenStudy (cookielate):

let me try to do it okay just give me a few minutes okay

OpenStudy (anonymous):

Yea i think i'll leave this one to cookie

OpenStudy (anonymous):

lol me too

OpenStudy (anonymous):

>-<

OpenStudy (cookielate):

do you have to answer one or write it for each step

OpenStudy (cookielate):

okay here is an example

OpenStudy (cookielate):

Step by step solution : Step 1 : Simplify x3+x2-4x - 4 Checking for a perfect cube : 1.1 x3+x2-4x-4 is not a perfect cube Trying to factor by pulling out : 1.2 Factoring: x3+x2-4x-4 Thoughtfully split the expression at hand into groups, each group having two terms : Group 1: x3+x2 Group 2: -4x-4 Pull out from each group separately : Group 1: (x+1) • (x2) Group 2: (x+1) • (-4) ------------------- Add up the two groups : (x+1) • (x^2-4) Which is the desired factorization Trying to factor as a Difference of Squares : Factoring: x^2-4 Theory : A difference of two perfect squares, A2−B2 can be factored into (A+B) • (A-B) Proof : (A+B) • (A-B) = A2−AB+BA−B62 = A2−AB+AB−B2 = A2 - B2 Note : AB = BA is the commutative property of multiplication. Note : - AB + AB equals zero and is therefore eliminated from the expression. Check : 4 is the square of 2 Check : x2 is the square of x1 Factorization is : (x + 2) • (x - 2) Final result : (x+2)∙(x−2)∙(x+1)

OpenStudy (anonymous):

ok sorry cookie but im still lost

OpenStudy (cookielate):

@starlord6200 are you suppose to pick an answer or solve each ?

OpenStudy (vickyisthesmartone123):

need help here?

OpenStudy (cookielate):

i gave him an example but he doesn't get it

OpenStudy (vickyisthesmartone123):

just give him the answer

OpenStudy (anonymous):

yes please and its multiple choice

OpenStudy (cookielate):

its not solving

OpenStudy (cookielate):

can u give me a minute @starlord6200 sorry if im taking long okay sorry :(

OpenStudy (anonymous):

its fine

OpenStudy (vickyisthesmartone123):

are you till working on this pellet?

OpenStudy (cookielate):

it not step 3

OpenStudy (cookielate):

its step 1

OpenStudy (cookielate):

for the first question

OpenStudy (anonymous):

no u have to figure out step 4 and 5

OpenStudy (cookielate):

for step 4 : (x2 + 9)(x + 1)

OpenStudy (cookielate):

step 5 (x + 3)(x + 3)(x + 1